# FLTrojan: Privacy Leakage Attacks against Federated Language Models through Selective Weight Tampering

### v3

@@ 5.5 Attack Cost Analysis
For VRI, storing the model weights at each iteration is not required. Only the candidates that pass the VRI test need to be stored. Additionally, the attacker needs to store only the weights of the selective sensitive layers of an LM instead of the entire model weights. For a Gemma-2b model, the total storage cost when we store the entire model weights ( 4.5 GB) in each iteration will be $O(N):=700*4.5=\textbf{3150 GB}$. But the number of layers in L* is 20 out of 164 total layers $(s=20/164)$, and if the fraction of rounds where victims participate is m=50%, the space complexity will be $O(m*s*N):=0.5*700*4.5*20/164=\textbf{~{}172 GB}$, significantly less than N. Additionally, quantization techniques are useful to reduce the space complexity of storing model checkpoints. By utilizing the Bits and Bytes  [^3] https://github.com/TimDettmers/bitsandbytes 8-bit quantization, we could reduce the stored model size by half ($\approx 2.2$ GB for the entire Gemma-2b model).
5.6 Benchmark Study
@@
Table 3: NSR comparison among our attacks with WTL, Decepticons [19] and three baselines on 200 canaries.
@@
We compare our results with Decepticons [19], which aims to uncover private user text by deploying malicious parameter vectors from the server. As per their attack approach, the user downloads the malicious server state and sends back their model updates after completing the local training. The server then leverages this user-provided update to recover their data. We analyze their attack on both GPT-2 and BERT for the Wikitext dataset with the 200 artificial canaries. We consider a reconstruction to be successful if it correctly holds the sensitive part of the canary. For GPT-2, it could reconstruct 137 canaries out of 200, whereas the NSR for BERT decreased to 42, as shown in Table 3. While their proposed method assumes the parameter server as an active adversary that sends malicious updates to the user, our static and dynamic attacks perform competitively with GPT-2 without any server cooperation. Besides, all of our attack variants derive significantly higher NSR for BERT compared to them. Finally, with the cooperation of an honest-but-curious server, our $dynamic^{+}$ attack marginally outperforms them with GPT-2, while the high NSR scores for FP and MU indicate that it is possible to cause greater privacy leakage if we consider the server as an active adversary. Besides, we also analyze the FILM attack [24] on GPT-2 which aims to recover private text data in federated learning. We put the related benchmark study for this work into the Appendix (H) due to space constraints.
@@
Figure 6: The tradeoff between model utility (in terms of perplexity) and private data memorization (in terms of exposure) with ($\epsilon=8$) and without DP
Metric No Defense $\mathbf{\epsilon=8}$ $\mathbf{\epsilon=2}$ $\mathbf{\epsilon=0.5}$ SCR
@@
Table 4: Number of Successful reconstructions (NSR) and model’s utility, i.e., average perplexity (Avg. PPX) in Enron email dataset with two defense mechanisms including Differential privacy with different privacy budgets ($\epsilon$) and Scrubbing (SCR)
5.7 Ablation Studies
5.7.1 Data Repetitions
Previous studies [39, 9] have shown that repeating sequences in the training set can increase memorization in language models. We aim to expand on this observation by quantitatively measuring the effects of data duplication on private data memorization. To do so, we repeat the canaries several times between 0 and 200 and add them to the local dataset of the victims. Our results, as shown in Figure 7(a), demonstrate a clear log-linear trend in memorization. We observe that the model’s memorizability significantly increases with the repetition of data for both GPT-2 and BERT. This is because the model tends to overfit on repeated data samples [75]. However, it is noticeable that the rate of increase in NSR quantity decreases for higher repetitions, i.e., both curves get less steeper. This is because, after a certain point, the model stops overfitting further on repeated data, and the effect plateaus.
5.7.2 Prompt Perturbation
To test the strength of our attack, we relax the assumption that the attacker has all of the private data prefixes in their original form. Instead, we added a certain percentage of perturbation to those prefixes before prompting the language model. To achieve this, we used the ChatGPT API [^4] https://openai.com/chatgpt to rephrase each of the prefixes by a specific percentage ($20\%-70\%$). Next, we used the BERT-embedding-based similarity score [37] to determine the semantic coherence between each rephrased prefix and its original counterpart. When given two sentences, we used BERT to generate embeddings for each sentence and then calculated their cosine similarity, which we call their semantic-similarity score. The results of this analysis are shown in Figure 7(b), where we display the number of successful reconstructions against different ranges of semantic-similarity scores. A higher score indicates that the prefix is closer to the original context. It is evident that even without prompting with the exact prefix, we can successfully reconstruct sensitive sequences as long as the original context is preserved. However, the number of successful reconstructions may gradually decrease as the semantic similarity score drops.
5.7.3 Number of Sensitive Layers
More ablation studies can be found in Appendix I
5.8 Differential Privacy
In machine learning, the DP SGD algorithm [1] is used to train neural networks with differentially private guarantees. To benchmark the performance of our attack against differential privacy, we implement differentially private (DP) training for the client language models [40]. Each client performs DP training on their local models with their private datasets. We train using various $\epsilon$ values and set $\delta=n^{-1.1}$ in each case, $n$ is the number of local training samples. The $\ell_{2}$-norm for gradient clipping is set to 0.01. More details about the algorithm are available in the Appendix C.
We conducted the experiment on the Enron email dataset and calculated exposure for all 125 private samples. We also picked 125 non-private samples from the victim’s local dataset and calculated the model’s perplexity for them. Figure 6 plots the exposure (EXP) of 125 private samples against the perplexity (PPX) of the corresponding 125 non-private samples. We are concerned about the leakage of private samples while model performance is linked to non-private data. From Figure 6, we see that DP reduces the EXP for the private samples. Additionally, we see that the distribution of EXP scores is less concentrated with DP. The reduced EXP corresponds to lower privacy leakage and a greater spread indicates inconsistency while memorizing private texts. The PPX increases with DP, indicating a reduction in utility. Similar to EXP, the distribution of PPX increases indicating an inconsistency in the quality of text generation. However, the EXP of other attack methods with DP is still higher than baseline LR. Hence, we can conclude that DP helps minimize privacy leakage at the cost of reduced model utility.
5.9 Scrubbing
@@
Scrubbing [10] is a method to identify and remove certain tokens from text. While it is frequently used for text anonymization [54], we use it for removing privacy-sensitive tokens from user text. Compared to canaries, it is more challenging to scrub out real-world private data, like Enron, since they are less structured and more stochastic. Nevertheless, we designed several regular expressions to capture the privacy-sensitive tokens in Enron and replace them with the token <UNK>. We managed to scrub out the sensitive tokens from almost 47% (58 out of 125) of the private sequences. As shown in Table 4, unlike DP, scrubbing reduces the NSR without hurting the model’s utility. However, applying DP and scrubbing further reduces NSR but greatly hurts the model’s utility.
5.10 Data Deduplication
@@
Data deduplication algorithms [39] remove duplicate training samples in large datasets before training language models. This process is used to reduce the amount of memorization in a trained language model [39, 7]. Examples that are repeated more often, as shown in Figure 7(a), are more likely to be extractable. Hence, removing multiple training sample occurrences will reduce the possibility of leakage to some extent.
5.11 Secure Aggregation
Secure aggregation protocols aim to solve the problem of computing a multiparty sum where none of the parties reveal their updates in the clear, even to the aggregator [5]. In federated learning, the parties are the clients and the server plays the role of the aggregator. In Equation 2, the clients contribute each individual $\theta^{i}$ in each epoch and all the parties securely aggregate the updates to only learn the summation $\theta$. Secure aggregation can be achieved by masking the individual updates using the shares of zero [5], homomorphic encryption [12], or functional encryption [74]. Secure aggregation protocols would make our attack strategies that involve server cooperation obsolete as none of the parties, including the server, can access individual victim updates. However, the remaining attacks can still be performed.
6 Future Work
Our proposed methods, which focus on language modeling tasks, could be extended to explore the impact of selective weight tampering in other domains, such as computer vision. Investigating how such tampering affects vision-based tasks could open new research directions. Since our study primarily deals with transformer models, a natural extension would be to apply these techniques to image transformers [57] in generative vision tasks. This could help assess whether similar vulnerabilities exist in image-based federated learning models and how privacy leakage could manifest in different modalities beyond language. Additionally, exploring hybrid models [15] or tasks involving multimodal data would further broaden the scope of these investigations.
7 Conclusion
@@
In this paper, we found that a malicious client in federated learning can confidently identify the rounds where the victims participate. By exploiting the model checkpoints from these rounds, the attacker can access more sensitive information than they could with just the final model. Furthermore, if the adversary manipulates the selective weights of the model using our proposed MDM techniques, privacy breaches can be intensified. Our most effective method increases the recall for membership inference by 29%, while also enabling up to 71% reconstruction of private data. However, current defenses like DP cannot defend against our attacks without significantly impacting the model’s utility. For future research, we recommend exploring more advanced federated learning frameworks and sophisticated defense mechanisms to guard against privacy breaches.
8 Ethical Considerations and Open Science
Federated learning is increasingly utilized in privacy-sensitive applications to train machine learning models on users’ data without explicitly accessing their local datasets. The FLTrojan attack poses a risk to users’ privacy as it actively increases memorization of private data and shows how data extraction and membership inference attacks can be launched on honest users participating in federated learning. We aim to shed light on the privacy pitfalls of federated learning and how existing defenses do not eliminate the threat of such attacks. Our attack methods highlight the importance of developing strong defenses to enhance the privacy of federated learning.
All datasets and source code associated with FLTrojan are publicly available [^5] Anonymous link: https://figshare.com/s/84afd0f40af3ac770684 to enhance the replicability and reproducibility of the results.
@@ Appendix A Federated Learning
Federated learning (FL) [46] proposes an algorithm for training machine learning models in a distributed setting. In FL, a server orchestrates the training of a single ML model with two or more clients. Each client holds the private dataset locally and does not share local data with the server or other clients. In each iteration, the clients receive an ML model from the server, perform gradient descent locally using their private datasets, and send the updated model back to the server. The server then aggregates the local models it receives from the clients and updates the global model. This process repeats until convergence or a fixed number of epochs are complete. The FL aggregation can be represented as follows:
@@ B.1 GPT-2
GPT-2 [59] is a transformer-based [69] neural language model developed by OpenAI. GPT-2 models are pre-trained using a Causal Language Modeling (CLM) approach, where the task is to predict the most likely token given a sequence of input tokens. Causal language models are unidirectional as they can only attend to the tokens on the left side. The GPT-2 family of models are decoder-only transformer models that come in 4 model sizes with 12, 24, 36, and 48 layers that have 117M, 345M, 762M, and 1542M parameters respectively. The model architecture is based on the original GPT [58] language model. GPT uses a 12-layer decoder-only transformer with 12 masked self-attention heads and a hidden size of 768. Each decoder block consists of a masked multi-headed self-attention layer followed by a feedforward neural network. Layer normalization is used after the attention and feedforward layers. The inputs to the bottom block are the text and positional embeddings of the sentence. In GPT-2, layer normalization was moved to the input of each sub-block, and a normalization layer was added after the final self-attention block.
B.2 BERT
@@
The BERT [13] family of models introduces bidirectional transformers. Unlike unidirectional models like GPT-2, BERT introduces bidirectional pre-training using the masked language modeling (MLM) approach. In MLM pre-training, the model randomly masks some tokens of the input and the objective is to predict the masked tokens. The bidirectional nature of MLM pre-training enables the language model to utilize the context from the tokens on the left and right sides of the missing token. The BERT models come in 2 sizes. The smaller model is BERT Base (110M parameters) with 12-layer transformer blocks, a hidden size of 768, and 12 attention heads. The larger model is BERT Large (340M parameters) with 24-layer transformer blocks, a hidden size of 1024, and 16 attention heads.
B.3 Gemma

### v4

@@ VI-B Attack Cost Analysis
For VRI, storing the model weights at each iteration is not required. Only the candidates that pass the VRI test need to be stored. Additionally, the attacker needs to store only the weights of the selective sensitive layers (L*) of an LM instead of the entire base weights. For a Gemma-2b model, with 4-bit QLoRA, the total storage cost when we store the entire model weights (~4.5 GB) in each iteration will be $O(N):=700*4.5/4=\textbf{787.5 GB}$. But the number of layers in L* is 20 out of 164 total layers $(s=20/164)$, and if the fraction of rounds where victims participate is m=50%, the space complexity will be $O(m*s*N):=0.5*787.5*20/164=\textbf{\textasciitilde 48 GB}$, which is ~93% less than $O(N)$, making it highly feasible for an attacker.
VI-C Benchmark Study
@@
TABLE III: NSR comparison among our attacks with WTL, Decepticons [fowl2022decepticons] and three baselines on 200 canaries.
@@
We compare our results with Decepticons [fowl2022decepticons], which aims to uncover private user text by deploying malicious parameter vectors from the server. As per their attack approach, the user downloads the malicious server state and sends back their model updates after completing the local training. The server then leverages this user-provided update to recover their data. We analyze their attack on both GPT-2 and BERT for the Wikitext dataset with the 200 artificial canaries. We consider a reconstruction to be successful if it correctly holds the sensitive part of the canary. For GPT-2, it could reconstruct 137 canaries out of 200, whereas the NSR for BERT decreased to 42, as shown in Table III. While their proposed method assumes the parameter server as an active adversary that sends malicious updates to the user, our static and dynamic attacks perform competitively with GPT-2 without any server cooperation. Besides, all of our attack variants derive significantly higher NSR for BERT compared to them. Finally, with the cooperation of an honest-but-curious server, our $dynamic^{+}$ attack marginally outperforms them with GPT-2, while the high NSR scores for FP and MU indicate that it is possible to cause greater privacy leakage if we consider the server as an active adversary. Besides, we also analyze the FILM attack [gupta2022recovering] on GPT-2 which aims to recover private text data in federated learning. We put the related benchmark study for this work into the Appendix (H) due to space constraints.
@@
Figure 6: The tradeoff between model utility (in terms of perplexity) and private data memorization (in terms of exposure) (a) with DP ($\epsilon=8$) (b) with DEPN and Add-DEPN
VI-D Privacy Defenses
VI-D1 Differential Privacy
In machine learning, the DP-SGD algorithm [Abadi_2016] is used to train neural networks with differentially private guarantees. To benchmark the performance of our attack against differential privacy (DP), we implement DP training for the client language models [li2022large]. Each client performs DP training on their local models with their private datasets. We train using various $\epsilon$ values and set $\delta=n^{-1.1}$ in each case, $n$ is the number of local training samples. The $\ell_{2}$-norm for gradient clipping is set to 0.01. More details about the algorithm are available in Appendix C.
We conducted the experiment on the Enron email dataset and calculated exposure for all 125 private samples. We also picked 125 non-private samples from the victim’s local dataset and calculated the model’s perplexity for them. Figure 6(a) plots the EXP of 125 private samples against the PPX of the corresponding 125 non-private samples. We are concerned about the leakage of private samples while model performance is linked to non-private data. From Figure 6(a), we see that DP reduces the EXP for the private samples. Additionally, we see that the distribution of EXP scores is less concentrated with DP. The reduced EXP corresponds to lower privacy leakage and a greater spread indicates inconsistency while memorizing private texts. The PPX increases with DP, indicating a reduction in utility. Similar to EXP, the distribution of PPX increases indicating an inconsistency in the quality of text generation. However, the EXP of other attack methods with DP is still higher than baseline LR. Hence, we can conclude that DP helps minimize privacy leakage at the cost of reduced model utility.
VI-D2 Pruning/Regularization
A complementary line of defense is to directly regularize those internal neurons that are most responsible for leaking private information. We therefore evaluate DEPN [wu2023depn], a neuron-level privacy protection framework that first identifies “privacy neurons” via gradient-based attribution and then edits their activations to suppress memorization of private sequences. Here, we utilized the same local dataset as the DP experiment (125 private, 125 non-private Enron data) for evaluation. For each private sample, we compute privacy attribution scores for all FFN neurons and aggregate them over examples to obtain a global ranking. We then select the top $z\in\{100,500,1000\}$ neurons as privacy neurons and apply DEPN’s editor to those neurons only. We quantify privacy leakage as the average exposure across all private samples, and model utility using validation perplexity on the non-private samples in Figure 6(b).
In the original DEPN editor, the activations of the selected privacy neurons are simply set to zero at inference time, effectively erasing their contribution to the logits. Consistent with DEPN’s analysis, we observe that increasing the number of edited neurons monotonically reduces exposure, as shown by the green line in Figure 6(b). However, in our federated Enron setting, this hard erasure behaves as an extremely aggressive regularizer, severely damaging utility despite its reduced exposure.
To obtain a softer form of regularization, we introduce an additive privacy neuron editor (Add-DEPN) inspired by model-editing methods such as ROME [meng2022locating]. Instead of zeroing, we apply a small additive correction, i.e., $h^{(l)}_{k}=h^{(l)}_{k}-\Delta$ , where $\Delta$ is chosen as a function of the privacy attribution score: $\Delta=\eta\cdot\mathrm{Att}\!\left(w^{(l)}_{k}\right)$. This means neurons with higher privacy contribution get slightly suppressed, not destroyed. As depicted by the orange line in Figure 6(b), exposure is reduced almost as much as with DEPN’s zeroing editor, but the increase in perplexity is substantially smaller. For moderate values (e.g., zz=500), additive editing yields a sizable reduction in NSR, as shown in Table IV, while keeping validation perplexity close to that of the baseline undefended model. We recommend utilizing Add-DEPN as a practical regularization mechanism at the client side in order to mitigate such private information leakage.
Metric Undefended $\mathbf{\epsilon=8}$ $\mathbf{\epsilon=2}$
Add-DEPN
$\mathbf{z=500}$
SCR
@@ VI-D2 Pruning/Regularization
TABLE IV: Number of Successful reconstructions and model’s utility, i.e., average perplexity in Enron email dataset with two defenses: Differential privacy with different privacy budgets ($\epsilon$), Scrubbing (SCR), and Add-DEPN
VI-D3 Scrubbing
@@
It is a method [chowdhury2021adversarial] to identify and remove certain tokens from text. While it is frequently used for text anonymization [pilan2022text], we use it for removing privacy-sensitive tokens from user text. Compared to canaries, it is more challenging to scrub out real-world private data, like Enron, since they are less structured and more stochastic. Nevertheless, we designed several regular expressions to capture the privacy-sensitive tokens in Enron and replace them with the token <UNK>. We managed to scrub out the sensitive tokens from almost 47% of the private sequences. As shown in Table IV, unlike DP, scrubbing reduces the NSR without hurting the model’s utility. However, applying DP and scrubbing together (SCR + $\epsilon=8$) further reduces NSR but greatly hurts the model’s utility. On the contrary, applying regularization, i.e., Add-DEPN along with scrubbing derives similar privacy gain with much better validation performance.
VI-D4 Data Deduplication and Secure Aggregation
@@
Data deduplication algorithms [lee-etal-2022-deduplicating] remove duplicate training samples in large datasets before training language models. This process is used to reduce the amount of memorization in a trained language model [lee-etal-2022-deduplicating, carlini2023quantifying]. Examples that are repeated more often, as discussed in Appendix I-A, are more likely to be extractable. Hence, removing multiple occurrences of training samples certainly reduces the possibility of leakage to some extent. Secure aggregation is another plausible defense which we discuss in Appendix J due to space constraints. Apart from that, there are a few gradient-based defenses [chang2024gradient], such as signSGD [bernstein2018signsgd], which do not apply against our attack, since FedAvg does not allow gradient sharing.
So, to summarize our discussion on the privacy defenses, we recommend scrubbing the local corpus with common PII patterns and deduplication of repeated samples as a data preprocessing step, and then applying modest regularization techniques like Add-DEPN on the local model to substantially reduce leakage propensity of an FL client’s privacy-sensitive data.
VII Limitations and Future Work
A few limitations in our study open up promising directions for future work. First, we focus on language modeling in English with medium-scale transformer LMs. It would be interesting to see if larger production models and multilingual settings exhibit different memorization and tampering behavior. Besides, we benchmark a generic suite of defenses, but co-designing new FL protocols that are intrinsically robust to such weight tampering would be more impactful. Beyond language modeling, an important direction is to study whether similar vulnerabilities arise in other modalities. In particular, applying our framework to image transformers [radford2021learning] and hybrid or multimodal models [dosovitskiy2020image] in federated vision and vision–language tasks could reveal how privacy leakage manifests when parameters jointly encode text and non-text signals.
VIII Conclusion
@@
In this paper, we found that a malicious client in federated learning can confidently identify the rounds where some other clients participate with privacy-sensitive data. Without requiring any gradient information, the attacker can retrieve more private data by exploiting the model checkpoints from these interim rounds than they could with just the final model. Furthermore, if the adversary manipulates the selective weights of the model using our proposed MDM techniques, privacy breaches can be intensified. Our most effective method increases the recall for membership inference by 29%, while also enabling up to 71% reconstruction of private data. However, choosing a balanced combination of defense strategies could mitigate the privacy risk of an FL client without significantly hurting utility.
Symbol Description
$G$ Large language model
$\mathcal{D}_{reg}$ Regular English sentences
$\mathcal{D}_{sen}$ Out-of-distribution texts
$L$ All layers
$L_{\star}$ Subset of $L$ that undergo a substantial transformation when fine-tuning $G$ with $\mathcal{D}_{sen}$
$_{\ominus} LL$$\setminus L_{\star}$
$W_{\star}$ Weights of $L_{\star}$
$\Delta W_{\star}$ Change in weights $W_{\star}$ after fine-tuning $G$ with $\mathcal{D}_{sen}$
$W_{\ominus}$ Weights of $L_{\ominus}$
$\Delta W_{\ominus}$ Change in weights $W_{\ominus}$ after fine-tuning $G$ with $\mathcal{D}_{sen}$
$G_{i}$ Global model snapshot in round $i$
$G_{r}$ Model achieved by fine-tuning $G$ with $\mathcal{D}_{reg}$
$G_{s}$ Model achieved by fine-tuning $G$ with $\mathcal{D}_{reg}\cup\mathcal{D}_{sen}$
$\delta_{i}$ Norm of weight difference for each of the $L_{\star}$ layers between $G$ and $G_{i}$
$\delta_{r}$ Norm of weight difference for each of the $L_{\star}$ layers between $G$ and $G_{r}$
$\delta_{s}$ Norm of weight difference for each of the $L_{\star}$ layers between $G$ and $G_{s}$
$\widehat{G_{i}}$ Victim round global snapshot
$\delta_{v}$ Distinction in $W_{\star}$ between $\widehat{G_{i}}$ and $G_{r}$
TABLE V: List of symbols and their descriptions
Figure 7: Norms of weight changes in MLP and self-attention layers of all 18 transformer blocks after fine-tuning Gemma with (a) regular English texts and (b) privacy-sensitive texts. The embedding and normalization layers are not shown.
Figure 8: Exposure scores for Different Attack Strategies along with the Baselines for (a)Wikitext with GPT-2, (b) Wikitext with BERT, (c)PTB with GPT-2, (d) PTB with BERT. (e)Enron with GPT-2 (f)Enron with BERT. All scores are calculated over 200 unique canaries.
Algorithm 3 Victim Round Identification (VRI)
2:Output: Whether $G_{i}$ is a $\widehat{G_{i}}$ instance or not.
3:Fine-tune $G$ with $\mathcal{D}_{reg}$ and obtain $G_{r}$
4:Fine-tune $G$ with $\mathcal{D}_{reg}\cup D_{sen}$ and obtain $G_{s}$
[…]
10:if $t_{r}$ rejects $H_{0}$ AND $t_{s}$ accepts $H_{0}$ then
11:  $G_{i}$ is a $\widehat{G_{i}}$ instance
12:else
13:  $G_{i}$ is not a $\widehat{G_{i}}$ instance
14:end if
@@ Appendix A Federated Learning
Federated learning (FL) proposes an algorithm for training machine learning models in a distributed setting. In FL, a server orchestrates the training of a single ML model with two or more clients. Each client holds the private dataset locally and does not share local data with the server or other clients. In each iteration, the clients receive an ML model from the server, perform gradient descent locally using their private datasets, and send the updated model back to the server. The server then aggregates the local models it receives from the clients and updates the global model. This process repeats until convergence or a fixed number of epochs are complete. The FL aggregation can be represented as follows:
@@ B-A GPT-2
GPT-2 [radford2019language] is a transformer-based [vaswani_etal] neural language model developed by OpenAI. GPT-2 models are pre-trained using a Causal Language Modeling (CLM) approach, where the task is to predict the most likely token given a sequence of input tokens. Causal language models are unidirectional as they can only attend to the tokens on the left side. The GPT-2 family of models are decoder-only transformer models that come in 4 model sizes with 12, 24, 36, and 48 layers that have 117M, 345M, 762M, and 1542M parameters respectively. The model architecture is based on the original GPT [radford2018improving] language model. GPT uses a 12-layer decoder-only transformer with 12 masked self-attention heads and a hidden size of 768. Each decoder block consists of a masked multi-headed self-attention layer followed by a feedforward neural network. Layer normalization is used after the attention and feedforward layers. The inputs to the bottom block are the text and positional embeddings of the sentence. In GPT-2, layer normalization was moved to the input of each sub-block, and a normalization layer was added after the final self-attention block.
B-B BERT
@@
The BERT [devlin-etal-2019-bert] family of models introduces bidirectional transformers. Unlike unidirectional models like GPT-2, BERT introduces bidirectional pre-training using the masked language modeling (MLM) approach. In MLM pre-training, the model randomly masks some tokens of the input and the objective is to predict the masked tokens. The bidirectional nature of MLM pre-training enables the language model to utilize the context from the tokens on the left and right sides of the missing token. The BERT models come in 2 sizes. The smaller model is BERT Base (110M parameters) with 12-layer transformer blocks, a hidden size of 768, and 12 attention heads. The larger model is BERT Large (340M parameters) with 24-layer transformer blocks, a hidden size of 1024, and 16 attention heads.
B-C Gemma